OpenStudy (anonymous):

Daniel's basic cell phone rate each month is $29.95. Add to that $5.95 for voice mail and $2.95 for text messaging. This past month Daniel spent an additional C dollars on long distance. His total bill was $62.35. How much did Daniel spend on long distance?

OpenStudy (anonymous):

29.95+5.95+2.95= 38.85 38.85 + c = 62.35 Subtract 38.85 from both sides of the equation. c = 62.35-38.85 combine like terms c= $23.50

OpenStudy (texaschic101):

I agree with @laugh1forever ....it looks good to me :)